# Multi-Focused Video Group Activities Hashing

### v2

@@
Abstract
With the explosive growth of video data in various complex scenarios, quickly retrieving group activities has become an urgent problem. However, many tasks can only retrieve videos focusing on an entire video, not the activity granularity. To solve this problem, we propose a new STVH (spatiotemporal interleaved video hashing) technique for the first time. Through a unified framework, the STVH simultaneously models individual object dynamics and group interactions, capturing the spatiotemporal evolution on both group visual features and positional features. Moreover, in real-life video retrieval scenarios, it may sometimes require activity features, while at other times, it may require visual features of objects. We then further propose a novel M-STVH (multi-focused spatiotemporal video hashing) as an enhanced version to handle this difficult task. The advanced method incorporates hierarchical feature integration through multi-focused representation learning, allowing the model to jointly focus on activity semantics features and object visual features. We conducted comparative experiments on publicly available datasets, and both STVH and M-STVH can achieve excellent results. The code is available in https://github.com/qzm1997/qzm1997_GAR.
keywords:
@@ Abstract
1 Introduction
Video group activity, which consists of interactions among multiple objects, is the core of video analysis because it can express high-level semantic information. For example, robbery activities can alert automatically for surveillance monitors ([33, 31]), kills in volleyball matches might be searched for score calculation ([16, 17]), and goals in football games are replayed as they are exciting moments ([29, 41]). With the explosive growth of video data in various complex scenarios, finding and retrieving activities quickly has become an urgent task.
@@ 1 Introduction
Figure 1: Our methods and existing methods.
Hash learning ([25, 36]) is a high-speed and efficient technique widely used in large-scale data retrieval, which can keep the distance between hash codes consistent with the distance between the original data. As shown in Fig. 1a, existing video hashing methods encode videos from a global perspective rather than activity-focused encoding. Some activity-focused methods (non-hash methods) can only perform categorization ([42, 47, 4]), which cannot satisfy the speed requirement when the data volume is large. As shown in Fig. 1b, these methods only perform simple fusion operations (e.g., add or concat) of visual and positional features, which are unable to effectively perform activity modeling. Furthermore, in real-life video retrieval scenarios, sometimes it may need to emphasize activities (i.e., activity-focused hash), while other times it may need to emphasize visual features of objects (i.e., visual-focused hash). Take football match videos as an example, we may retrieve offensive segments that emphasize activities, or we may retrieve segments only from one team that emphasize visual features. Therefore, if only one set of hash codes can meet the above requirements, it can significantly reduce storage cost.
@@ 2.1 Video Hashing
Video hash retrieval methods have predominantly based on deep learning, which can be classified into supervised and unsupervised hash learning approaches. Supervised hash learning generally extracts video features via a neural network, followed by a hash code generation method through label-based supervision. [30] introduced a similarity-preserving deep temporal hashing network that unifies video modeling and hash learning into a single, cohesive process. [2] employed a 3D convolution ([14]) extracted spatiotemporal features from videos, subsequently mapping the extracted features into a unified binary space. [39] enabled an uncertainty-aware video hashing by predicting the probability distribution of hash codes, thus offering robust uncertainty quantification.
@@
Unsupervised video hash learning methods initially relied on autoencoders or clustering-based approaches for model training. [32] designed a hierarchical binary autoencoder, where a convolutional neural network serves as the video encoder and an LSTM ([12]) acts as the decoder, allowing the model to learn video features across multiple dimensions. [21] employed a transformer-based autoencoding network combined with time-sensitivity regularization, effectively minimizing sensitivity to local temporal disturbances while retaining global temporal sequence information. By utilizing a hash-based affinity matrix, this approach effectively preserves pairwise similarity between video samples. Recently, with the advent of contrastive learning techniques, some researchers have extended this unsupervised learning paradigm to video hash learning. [38] generated two distinct views of the video features through two different random masking techniques, then leverages contrastive learning to maximize the similarity of the hash codes between the views, followed by hash learning via video reconstruction from the hash codes. [19] built a re-ranking framework based on a knowledge distillation scheme and a selection mechanism that allows large unlabeled datasets to train our network of students and selectors. It not only improves the efficiency of retrieval but also ensures competitive performance.
2.2 Group Activity Recognition
@@
Group Activity recognition methods can be categorized into traditional and deep learning-based approaches, depending on the use of deep learning techniques. Traditional group behavior recognition predominantly depends on handcrafted feature extraction (e.g., [7], [26]), followed by the use of probabilistic graphical models like POMDPs or AND-OR logic for behavior prediction. Recently, with the rapid advancement of deep learning, architectures such as [15], [12], [18], and [6] have made significant strides in the field of group behavior recognition. Several state-of-the-art group activity recognition methods are predominantly based on deep neural network models. For instance, [46] organically combined joint motion, trajectories, and object positions to generate richer activity representations, further enhancing the corresponding features via gating mechanisms and self-attention mechanisms, leading to the final group behavior classification. [42] introduced a third-order active factor graph network that simulates the third-order interactions between every triplet of active object. To enhance group consistency, it incorporates a consistency-aware inference module, which includes two penalty terms that model the inconsistency between object and group activity. [10] investigated spatial and temporal correlations, proposing a novel loss function for self-supervised group action recognition. [17] introduced a random walk graph convolutional network for group activity recognition, incorporating a Levy flight random walk mechanism within GCN to capture information from various nodes, while leveraging prior positional information to recognize group activities. [23] utilized CNN to extract image features and then design an action-centered aggregation strategy, grouping objects performing different actions before making predictions. [45] introduced a spatiotemporal dynamic reasoning model that predicts the relational matrix and captures dynamic walking offsets through a joint processing method that integrates dynamic relation and dynamic walking modules. [47] presented a multi-level spatiotemporal Transformer-based relational reasoning framework aimed at exploring temporal dependencies and spatial dynamics among different objects in group activities.
2.3 Summary
@@
Figure 2: STVH consists of four main modules: 1) Visual Module: extracting features from the input video; 2) Positional Module: modeling spatiotemporal relationships on the input position information; 3) Spatiotemporal Interleaving Module: interleaving visual features and positional features; and 4) Hashing and Classification Learning Module: fusing the visual features with the position features and outputting the corresponding hash codes and classifications.
3 Problem Definition
Assume the training set input video is $\boldsymbol{X}=\{\boldsymbol{x}_{1},\boldsymbol{x}_{2},...,\boldsymbol{x}_{M}\}$ and the position information is $\boldsymbol{Boxes}=\{\boldsymbol{boxes}_{1},\boldsymbol{boxes}_{2},...,\boldsymbol{boxes}_{M}\}$, which has a activity class label of $\boldsymbol{Y}^{acty}=\{\boldsymbol{y}^{acty}_{1},\boldsymbol{y}^{acty}_{2},...,\boldsymbol{y}^{acty}_{M}\}$, where $M$ denotes the number of video samples. $\boldsymbol{x}_{i}\in\mathbb{R}^{T\times C\times W\times H}$ denotes that the input video has $T$ frames, and the scale of each frame is $C\times W\times H$. $\boldsymbol{boxes}_{i}\in\mathbb{R}^{T\times N\times 4}$ is the position information of $N$ objects in video $\boldsymbol{x}_{i}$, and $\boldsymbol{box}_{j_{t}}$ is the position information of the $j$-th object in video $\boldsymbol{x}_{i}$ on the $t$-th frame. $\boldsymbol{Y}^{action}=\{\boldsymbol{y}^{action}_{i1},\boldsymbol{y}^{action}_{i2},...,\boldsymbol{y}^{action}_{iN}\}$ denotes the object action in video $\boldsymbol{x}_{i}$. The objective of STVH is to map the input video and location information into a $K$-bits hash code while M-STVH is to map these inputs into multiple $K$-bits hash code with a different focus. The specific notation description is shown in Table 1.
Table 1: Explanatory Table of Selected Symbols
@@ 4 STVH Model
4.1 Visual Module
The vision module extracts visual features from each object in the input video $\boldsymbol{x}_{i}$. First, we perform multiscale features extraction on the video frames using an ImageNet-pretrained Inception-v3 model ([34]), yielding a feature tensor $\boldsymbol{f}\in\mathbb{R}^{T\times d_{v}\times W^{\prime}\times H^{\prime}}$} where $d_{v}=512$ and $W^{\prime}$ and $H^{\prime}$ are the scales of the feature map after downsampling $W$ and $H$ 32 times. We then extract object visual features $\boldsymbol{f}_{roi}\in\mathbb{R}^{N\times T\times d_{v}\times W^{\prime\prime}\times H^{\prime\prime}}$ by their corresponding positional $\boldsymbol{boxes}_{i}$ via RoIAlign, where $W^{\prime\prime}$ and $H^{\prime\prime}$ are both 5. To compactly represent these features, we apply sequential 3D and 2D convolutions to vectorize $\boldsymbol{f}_{roi}$, producing a unified visual feature $\boldsymbol{f}_{v}\in\mathbb{R}^{N\times T\times d}$} with a dimension of $d=1024$.
4.2 Positional Module
Although visual features are important in inferring group activities, relying solely on object features still poses challenges. For example, “running" and “jogging" may exhibit nearly identical visual patterns, making them difficult to distinguish without additional cues. To address this, we analyze the spatiotemporal positional changes of objects, which provide critical discriminative signals for activity recognition. Specifically, our module takes the positional coordinates of all objects in the video as input and models their trajectories to capture activity-specific motion dynamics.
Figure 3: Modeling the action based on the computational IoU.
@@ 4.2 Positional Module
$\operatorname{IoU}\left(j,t_{1},t_{2}\right)=\frac{\boldsymbol{box}_{j_{t_{1}}}\cap\boldsymbol{box}_{j_{t_{2}}}}{\boldsymbol{box}_{j_{t_{1}}}\cup\boldsymbol{box}_{j_{t_{2}}}},$ (1)
@@
$\boldsymbol{dist}_{ij}=1-\sqrt{\frac{\left(\boldsymbol{box}_{i}-\boldsymbol{box}_{j}\right)^{2}}{\boldsymbol{std}}},$ (2)
@@
4.3 Spatiotemporal Interleaving Module
@@
Figure 4: Sparse graph relation attention module (SGAT), fusing visual features as well as positional features at an attention.
@@
$\begin{array}[]{l}\boldsymbol{f}^{\prime}_{(\tau)}=\text{LayerNorm}(\boldsymbol{f}_{(\tau)}),\\ \\ \boldsymbol{f}^{\prime\prime}_{(\tau)}=\text{SGAT}(\boldsymbol{f}^{\prime}_{(\tau)},\boldsymbol{G}),\\ \\ \boldsymbol{f}_{(\tau+1)}=\text{FNN}(\boldsymbol{f}^{\prime\prime}_{(\tau)}),\end{array}$ (3)
@@
$\begin{array}[]{l}\operatorname{SGAT}(\boldsymbol{f},\boldsymbol{G})=\boldsymbol{W}_{1}\boldsymbol{f}\text{ Attention }(\boldsymbol{f},\boldsymbol{G}),\\ \\ \text{ Attention }(\boldsymbol{f},\boldsymbol{G})=\operatorname{AT}_{v}(\boldsymbol{f})\times\operatorname{AT}_{p}(\boldsymbol{G}),\\ \\ \operatorname{AT}_{v}(\boldsymbol{f})=\operatorname{SoftMax}\left(\frac{\left(\boldsymbol{W}_{2}\boldsymbol{f}\right)\left(\boldsymbol{W}_{3}\boldsymbol{f}\right)^{T}}{\sqrt{C}}\right),\\ \\ \operatorname{AT}_{p}(\boldsymbol{G})=\operatorname{SoftMax}(\boldsymbol{G}\times\boldsymbol{W}),\end{array}$ (4)
@@
4.4 Hashing and Classification Learning Module
@@
$\begin{array}[]{l}\boldsymbol{h}_{i}=\text{sign}(\text{AVG}(\boldsymbol{F}_{h}(\boldsymbol{f}_{S}))),\\ \\ \widetilde{\boldsymbol{y}}_{i}^{acty}=\boldsymbol{F}_{c}(\boldsymbol{h}_{i}),\end{array}$ (5)
4.5 Loss Function
@@
$\begin{array}[]{c}L_{\text{acty }}=\sum_{i=1}^{M}-\boldsymbol{y}_{i}^{\text{acty }}\log\left(\widetilde{\boldsymbol{y}}_{i}^{{acty}}\right),\\ L_{\text{action }}=\sum_{i=1}^{M}\sum_{j=1}^{N}-\boldsymbol{y}_{ij}^{\text{action }}\log\left(\widetilde{\boldsymbol{y}}_{ij}^{{action}}\right).\end{array}$ (6)
@@
$L_{q}=\sum_{i=1}^{M}\sum_{j=1}^{M}exp(\frac{1}{K}|\boldsymbol{h}_{i}^{T}\boldsymbol{h}_{j}-\boldsymbol{b}_{i}^{T}\boldsymbol{b}_{j}|),$ (7)
@@
$\begin{array}[]{c}L_{con(i,i)}=\sum_{i=1}^{B}\log\frac{\sum_{j=1}^{B}\left({sim}\left(\boldsymbol{a}_{i},\boldsymbol{b}_{j}\right)\right)+\left({sim}\left(\boldsymbol{a}_{j},\boldsymbol{b}_{i}\right)\right)}{{sim}\left(\boldsymbol{a}_{i},\boldsymbol{b}_{i}\right)},\\ \\ {sim}\left(\boldsymbol{a}_{i},\boldsymbol{b}_{{i}}\right)=\exp(\frac{\boldsymbol{a}_{i}^{T}\boldsymbol{b}_{i}}{\left\|\boldsymbol{a}_{i}\right\|_{2}\left\|\boldsymbol{b}_{i}\right\|_{2}}),\\ \\ L_{CON}=\sum_{i=1}^{B}L_{con(i,i)},\end{array}$ (8)
4.6 Optimization

### v3

@@
Abstract.
With the explosive growth of video data in various complex scenarios, quickly retrieving group activities has become an urgent problem. However, many tasks can only retrieve videos focusing on an entire video, not the activity granularity. To solve this problem, we propose a new STVH (spatiotemporal interleaved video hashing) technique for the first time. Through a unified framework, the STVH simultaneously models individual object dynamics and group interactions, capturing the spatiotemporal evolution on both group visual features and positional features. Moreover, in real-life video retrieval scenarios, it may sometimes require activity features, while at other times, it may require visual features of objects. We then further propose a novel M-STVH (multi-focused spatiotemporal video hashing) as an enhanced version to handle this difficult task. The advanced method incorporates hierarchical feature integration through multi-focused representation learning, allowing the model to jointly focus on activity semantics features and object visual features. We conducted comparative experiments on publicly available datasets, and both STVH and M-STVH can achieve excellent results.
@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: TOISccs: Information systems  Top-k retrieval in databases
1. Introduction
Video group activity, which consists of interactions among multiple objects, is the core of video analysis because it can express high-level semantic information. For example, robbery activities can alert automatically for surveillance monitors (real1; Social), kills in volleyball matches might be searched for score calculation (VD; RWGCN), and goals in football games are replayed as they are exciting moments (soccer1; soccer2). With the explosive growth of video data in various complex scenarios, finding and retrieving activities quickly has become an urgent task.
@@ 1. Introduction
Figure 1. Our methods and existing methods.
Hash learning (hash1; hash2; hashing; image_hash) is a high-speed and efficient technique widely used in large-scale data retrieval, which can keep the distance between hash codes consistent with the distance between the original data. As shown in Fig. 1a, existing video hashing methods encode videos from a global perspective rather than activity-focused encoding. Some activity-focused methods (non-hash methods) can only perform categorization (AFGf; MLST; CAD), which cannot satisfy the speed requirement when the data volume is large. As shown in Fig. 1b, these methods only perform simple fusion operations (e.g., add or concat) of visual and positional features, which are unable to effectively perform activity modeling. Furthermore, in real-life video retrieval scenarios, sometimes it may need to emphasize activities (i.e., activity-focused hash), while other times it may need to emphasize visual features of objects (i.e., visual-focused hash). Take football match videos as an example, we may retrieve offensive segments that emphasize activities, or we may retrieve segments only from one team that emphasize visual features. Therefore, if only one set of hash codes can meet the above requirements, it can significantly reduce storage cost.
@@ 2.1. Video Hashing
Video hash retrieval methods have predominantly based on deep learning, which can be classified into supervised and unsupervised hash learning approaches. Supervised hash learning generally extracts video features via a neural network, followed by a hash code generation method through label-based supervision. SPTDH (SPTDH) introduced a similarity-preserving deep temporal hashing network that unifies video modeling and hash learning into a single, cohesive process. DVSH (DVSH) employed a 3D convolution(3D) extracted spatiotemporal features from videos, subsequently mapping the extracted features into a unified binary space. BrVAE (BrVAE) enabled an uncertainty-aware video hashing by predicting the probability distribution of hash codes, thus offering robust uncertainty quantification.
@@
Unsupervised video hash learning methods initially relied on autoencoders or clustering-based approaches for model training. SSVH (SSVH) designed a hierarchical binary autoencoder, where a convolutional neural network serves as the video encoder and an LSTM (LSTM) acts as the decoder, allowing the model to learn video features across multiple dimensions. TSVH (TSVH) employed a transformer-based autoencoding network combined with time-sensitivity regularization, effectively minimizing sensitivity to local temporal disturbances while retaining global temporal sequence information. By utilizing a hash-based affinity matrix, this approach effectively preserves pairwise similarity between video samples. Recently, with the advent of contrastive learning techniques, some researchers have extended this unsupervised learning paradigm to video hash learning. COMH (COMH) generated two distinct views of the video features through two different random masking techniques, then leverages contrastive learning to maximize the similarity of the hash codes between the views, followed by hash learning via video reconstruction from the hash codes. Dns (Dns) built a re-ranking framework based on a knowledge distillation scheme and a selection mechanism that allows large unlabeled datasets to train our network of students and selectors. It not only improves the efficiency of retrieval but also ensures competitive performance. KPSC-P and KPSC-F (KPSC) utilized pre-trained visual language models as knowledge sources while prioritizing video functionality to enhance diversity and reduce noise, ultimately achieving unsupervised video retrieval.
2.2. Group Activity Recognition
@@
Group Activity recognition methods can be categorized into traditional and deep learning-based approaches, depending on the use of deep learning techniques. Traditional group behavior recognition predominantly depends on handcrafted feature extraction (e.g., HOG (HOG), SIFT (SIFT)), followed by the use of probabilistic graphical models like POMDPs or AND-OR logic for behavior prediction. Recently, with the rapid advancement of deep learning, architectures such as ResNet (ResNet), LSTM (LSTM), GCN (GCN), and GRU (GRU) have made significant strides in the field of group behavior recognition. Several state-of-the-art group activity recognition methods are predominantly based on deep neural network models. For instance, MOGAR (MOGAR) organically combined joint motion, trajectories, and object positions to generate richer activity representations, further enhancing the corresponding features via gating mechanisms and self-attention mechanisms, leading to the final group behavior classification. AFGNet (AFGf) introduced a third-order active factor graph network that simulates the third-order interactions between every triplet of active object. To enhance group consistency, it incorporates a consistency-aware inference module, which includes two penalty terms that model the inconsistency between object and group activity. (STAHs) investigated spatial and temporal correlations, proposing a novel loss function for self-supervised group action recognition. RWGCN (RWGCN) introduced a random walk graph convolutional network for group activity recognition, incorporating a Levy flight random walk mechanism within GCN to capture information from various nodes, while leveraging prior positional information to recognize group activities. ASTFormer (ASTFormer) utilized CNN to extract image features and then design an action-centered aggregation strategy, grouping objects performing different actions before making predictions. DIN (DIN) introduced a spatiotemporal dynamic reasoning model that predicts the relational matrix and captures dynamic walking offsets through a joint processing method that integrates dynamic relation and dynamic walking modules. MLST-Former (MLST) presented a multi-level spatiotemporal Transformer-based relational reasoning framework aimed at exploring temporal dependencies and spatial dynamics among different objects in group activities.
2.3. Summary
@@
Figure 2. STVH consists of four main modules: 1) Visual Module: extracting features from the input video; 2) Positional Module: modeling spatiotemporal relationships on the input position information; 3) Spatiotemporal Interleaving Module: interleaving visual features and positional features; and 4) Hashing and Classification Learning Module: fusing the visual features with the position features and outputting the corresponding hash codes and classifications.
3. Problem Definition
Assume the training set input video is $\boldsymbol{X}=\{\boldsymbol{x}_{1},\boldsymbol{x}_{2},...,\boldsymbol{x}_{M}\}$ and the position information is $\boldsymbol{Boxes}=\{\boldsymbol{boxes}_{1},\boldsymbol{boxes}_{2},\\ ...,\boldsymbol{boxes}_{M}\}$, which has a activity class label of $\boldsymbol{Y}^{acty}=\{\boldsymbol{y}^{acty}_{1},\boldsymbol{y}^{acty}_{2},...,\boldsymbol{y}^{acty}_{M}\}$, where $M$ denotes the number of video samples. $\boldsymbol{x}_{i}\in\mathbb{R}^{T\times C\times W\times H}$ denotes that the input video has $T$ frames, and the scale of each frame is $C\times W\times H$. $\boldsymbol{boxes}_{i}\in\mathbb{R}^{T\times N\times 4}$ is the position information of $N$ objects in video $\boldsymbol{x}_{i}$, and $\boldsymbol{box}_{j_{t}}$ is the position information of the $j$-th object in video $\boldsymbol{x}_{i}$ on the $t$-th frame. $\boldsymbol{Y}^{action}=\{\boldsymbol{y}^{action}_{i1},\boldsymbol{y}^{action}_{i2},...,\boldsymbol{y}^{action}_{iN}\}$ denotes the object action in video $\boldsymbol{x}_{i}$. The objective of STVH is to map the input video and location information into a $K$-bits hash code while M-STVH is to map these inputs into multiple $K$-bits hash code with a different focus. The specific notation description is shown in Table 1.
Table 1. Explanatory Table of Selected Symbols
@@ 4. STVH Model
4.1. Visual Module
The vision module extracts visual features from each object in the input video $\boldsymbol{x}_{i}$. First, we perform multiscale features extraction on the video frames using an ImageNet-pretrained Inception-v3 model ((inv3)), yielding a feature tensor $\boldsymbol{f}\in\mathbb{R}^{T\times d_{v}\times W^{\prime}\times H^{\prime}}$} where $d_{v}=512$ and $W^{\prime}$ and $H^{\prime}$ are the scales of the feature map after downsampling $W$ and $H$ 32 times. We then extract object visual features $\boldsymbol{f}_{roi}\in\mathbb{R}^{N\times T\times d_{v}\times W^{\prime\prime}\times H^{\prime\prime}}$ by their corresponding positional $\boldsymbol{boxes}_{i}$ via RoIAlign, where $W^{\prime\prime}$ and $H^{\prime\prime}$ are both 5. To compactly represent these features, we apply sequential 3D and 2D convolutions to vectorize $\boldsymbol{f}_{roi}$, producing a unified visual feature $\boldsymbol{f}_{v}\in\mathbb{R}^{N\times T\times d}$} with a dimension of $d=1024$.
4.2. Positional Module
Although visual features are important in inferring group activities, relying solely on object features still poses challenges. For example, “running” and “jogging” may exhibit nearly identical visual patterns, making them difficult to distinguish without additional cues. To address this, we analyze the spatiotemporal positional changes of objects, which provide critical discriminative signals for activity recognition. Specifically, our module takes the positional coordinates of all objects in the video as input and models their trajectories to capture activity-specific motion dynamics.
Figure 3. Modeling the action based on the computational IoU.
@@ 4.2. Positional Module
(1) $\operatorname{IoU}\left(j,t_{1},t_{2}\right)=\frac{\boldsymbol{box}_{j_{t_{1}}}\cap\boldsymbol{box}_{j_{t_{2}}}}{\boldsymbol{box}_{j_{t_{1}}}\cup\boldsymbol{box}_{j_{t_{2}}}},$
@@
(2) $\boldsymbol{dist}_{ij}=1-\sqrt{\frac{\left(\boldsymbol{box}_{i}-\boldsymbol{box}_{j}\right)^{2}}{\boldsymbol{std}}},$
@@
4.3. Spatiotemporal Interleaving Module
@@
Figure 4. Sparse graph relation attention module (SGAT), fusing visual features as well as positional features at an attention.
@@
(3) $\begin{array}[]{l}\boldsymbol{f}^{\prime}_{(\tau)}=\text{LayerNorm}(\boldsymbol{f}_{(\tau)}),\\ \\ \boldsymbol{f}^{\prime\prime}_{(\tau)}=\text{SGAT}(\boldsymbol{f}^{\prime}_{(\tau)},\boldsymbol{G}),\\ \\ \boldsymbol{f}_{(\tau+1)}=\text{FNN}(\boldsymbol{f}^{\prime\prime}_{(\tau)}),\end{array}$
@@
(4) $\begin{array}[]{l}\operatorname{SGAT}(\boldsymbol{f},\boldsymbol{G})=\boldsymbol{W}_{1}\boldsymbol{f}\text{ Attention }(\boldsymbol{f},\boldsymbol{G}),\\ \\ \text{ Attention }(\boldsymbol{f},\boldsymbol{G})=\operatorname{AT}_{v}(\boldsymbol{f})\times\operatorname{AT}_{p}(\boldsymbol{G}),\\ \\ \operatorname{AT}_{v}(\boldsymbol{f})=\operatorname{SoftMax}\left(\frac{\left(\boldsymbol{W}_{2}\boldsymbol{f}\right)\left(\boldsymbol{W}_{3}\boldsymbol{f}\right)^{T}}{\sqrt{C}}\right),\\ \\ \operatorname{AT}_{p}(\boldsymbol{G})=\operatorname{SoftMax}(\boldsymbol{G}\times\boldsymbol{W}),\end{array}$
@@
4.4. Hashing and Classification Learning Module
@@
(5) $\begin{array}[]{l}\boldsymbol{h}_{i}=\text{sign}(\text{AVG}(\boldsymbol{F}_{h}(\boldsymbol{f}_{S}))),\\ \\ \widetilde{\boldsymbol{y}}_{i}^{acty}=\boldsymbol{F}_{c}(\boldsymbol{h}_{i}),\end{array}$
4.5. Loss Function
@@
(6) $\begin{array}[]{c}L_{\text{acty }}=\sum_{i=1}^{M}-\boldsymbol{y}_{i}^{\text{acty }}\log\left(\widetilde{\boldsymbol{y}}_{i}^{{acty}}\right),\\ L_{\text{action }}=\sum_{i=1}^{M}\sum_{j=1}^{N}-\boldsymbol{y}_{ij}^{\text{action }}\log\left(\widetilde{\boldsymbol{y}}_{ij}^{{action}}\right).\end{array}$
@@
(7) $L_{q}=\sum_{i=1}^{M}\sum_{j=1}^{M}exp(\frac{1}{K}|\boldsymbol{h}_{i}^{T}\boldsymbol{h}_{j}-\boldsymbol{b}_{i}^{T}\boldsymbol{b}_{j}|),$
@@
(8) $\begin{array}[]{c}L_{con(i,i)}=\sum_{i=1}^{B}\log\frac{\sum_{j=1}^{B}\left({sim}\left(\boldsymbol{a}_{i},\boldsymbol{b}_{j}\right)\right)+\left({sim}\left(\boldsymbol{a}_{j},\boldsymbol{b}_{i}\right)\right)}{{sim}\left(\boldsymbol{a}_{i},\boldsymbol{b}_{i}\right)},\\ \\ {sim}\left(\boldsymbol{a}_{i},\boldsymbol{b}_{{i}}\right)=\exp(\frac{\boldsymbol{a}_{i}^{T}\boldsymbol{b}_{i}}{\left\|\boldsymbol{a}_{i}\right\|_{2}\left\|\boldsymbol{b}_{i}\right\|_{2}}),\\ \\ L_{CON}=\sum_{i=1}^{B}L_{con(i,i)},\end{array}$
4.6. Optimization